# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v1

@@ 9.3 What kind of learning capability should a general manipulation framework possess?
Training RFMs simultaneously requires a large amount of training data. Current data collection methods include gathering data in simulation, using skill library to add data, teleoperation, and imitation learning from human videos. While collecting data in simulation is cost-effective, it suffers from the sim-to-real gap. In order to add data using skill library, it is necessary to construct an initial skill library. Teleoperation can address the sim-to-real gap but comes with high hardware costs and the issue of unnatural third-person demonstration data. Imitation learning from human videos involves mapping human trajectories from videos to robot space, leading to issues with trajectory accuracy. This paper argues that the optimal data collection method involves individuals wearing exoskeleton devices equipped with various sensors. These devices accurately capture human actions and map them to robot operation space, enabling first-person data collection and addressing the sim-to-real gap.
@@ 9.5 How to establish a benchmark?
The challenge in establishing benchmarks for foundation models for manipulation lies in addressing diverse task objectives. Thus, a framework categorizing benchmarks for different tasks is necessary. This paper’s framework encompasses multiple tasks including interaction, hierarchical, perception, pre- and post-conditions detection, policy, and manipulation data generation. VLADP models cannot achieve transfer ability across different robot configurations, and the workspaces of different robot configurations also differ. Designing an evaluation function to incorporate robot configurations may alleviate this issue. Moreover, the real-time performance of foundation models also needs consideration.
@@ 10 Conclusion
In this survey, we aim to outline the opportunities brought by foundation models for a general-purpose robot framework. Except for RFMs, none of the foundation models have utilized manipulation task-level data to train, thus only capable of addressing a subset of functionalities for general operations. The current framework for RFMs resembles human learning, requiring no calibration or ontology parameters, directly learning the mapping between vision and action. However, this demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring safety remains an unresolved concern.
As we propose a framework for a general-purpose manipulation agent, we first consider it to be language-conditioned, hence integrating an interaction module. The ability to understand corrective instruction within this module is crucial; agents inevitably make mistakes, and being able to correct them is pivotal on top of the three robotic fundamental principles. Understanding the environment and task execution in Pre- and Post-conditions Detection determines how tasks are completed and their measurement, while Post-hoc Correction serves as a check and self-correcting mechanism for robot safety and smooth task execution. Optimizing task execution plans within the Hierarchy of Skills is a decision-making process determining the ’intelligence’ of this agent, while skill execution decides the ’physical fitness’ of the agent.
Foundation models accelerate the research progress of general agents. LLMs and VFMs facilitate open-world interaction and perception, VLMs and LMMs ensure multimodal understanding. Current multimodal large models align modalities to text, as text tokens are conducive to reasoning, akin to humans contemplating information from books for cognitive processes. VGMs present significant opportunities for scene or object generation. However, each foundation model enhances the capability of zero-shot and few-shot learning. Autonomous driving is classified into levels L0-L5 . If we rank improving old skills and rigid static short-horizon operations as L0, the development of foundation models will bring excellent opportunities for general manipulation at L0.
While this survey raises some questions for consideration, the proposed framework for a general-purpose robot is sequential, whereas human behavior is often parallel. Nonetheless, we consider a sequentially connected framework as the first step towards L0 general-purpose manipulation agent. However, it’s undeniable that each module within the current framework faces numerous challenges:
@@
    Interaction In human interaction, there’s more than just language—it involves gestures or actions too. Adding multiple modal inputs to interaction modules can enhance recognition capabilities.
@@
    Hierarchical of skills The hierarchy of skills still has many unconsidered factors, such as achieving tasks in the shortest time with the highest efficiency, and how to devise proactive strategies for dynamic scenes.
@@
    Pre- and post-conditions detection Post-conditions detection needs to analyze reasons for task failure and enables self-correction within the post-hoc correction module. In current research, failure in task execution is typically determined by analyzing the status of objects after skill execution. But, post-conditions detection should instead analyze the robot’s own execution process to identify reasons for task failure. For instance, if a robot fails a task due to object collision during execution, post-hoc correction should prevent such occurrences during corrective actions. The perspective of observation might be a hidden challenge, requiring agents to actively perceive to analyze failure reason.
@@
    State The representation of state requires integration of multiple modalities, such as touch and hearing. At the same time, it’s important to consider the opportunities that foundation models can bring to active perception.
@@
    Policy The policy encompasses various RFMs, yet RFMs lack the capacity for self-exploration and a unified benchmark for comparing them, hindering RFMs development. Meanwhile, the extensive parameters of RFMs require significant computational resources for real-time operation, and model training also demands abundant data.
@@
    Environment Transition Module The foundation models inherently contain abundant physical priors. Applying foundation models to build a highly realistic physical model assist reinforcement learning training in transition models is a direction worth exploring.
@@
    Data Generation Output from LLMs and VGMs in data collection remains unstable, necessitating appropriate check module and data cleaning algorithms.

### v2

@@ 9.3 What kind of learning capability should a general manipulation framework possess?
Training RFMs simultaneously requires a large amount of training data. Current data collection methods include gathering data in simulation, using skill library to generate data, teleoperation, and imitation learning from human videos. While collecting data in simulation is cost-effective, it suffers from the sim-to-real gap. In order to generate data using skill library, it is necessary to construct an initial skill library. Teleoperation can address the sim-to-real gap but comes with high hardware costs and the issue of unnatural third-person demonstration data. Imitation learning from human videos involves mapping human trajectories from videos to robot space, leading to issues with trajectory accuracy. This paper argues that the optimal data collection method involves individuals wearing exoskeleton device equipped with various sensors. This device accurately capture human actions and map the actions to robot space, enabling first-person data collection and addressing the sim-to-real gap.
9.5 How to enable general manipulation with dexterous capability?
Currently, foundation models for manipulation focus mainly on simple tasks like ”Pushing, Pulling, Grasping, and Placing.” Complex tasks like ”Reorientation & Relocation, Screwing, and In-hand Manipulation” have not been widely studied. The primary difference between these complex tasks and simple tasks is that these complex tasks involve dynamic contact points, high real-time requirements, and contact-rich. The primary challenge lies in the data quality and the complexity of the training process. To address this issue, it is necessary to enhance the Skill Execution module of the framework presented in this paper, with a particular focus on the Policy. However, data and training methods for dexterous manipulation are closely tied to the challenges faced in dexterous manipulation. Therefore, this section explores how to enable general manipulation with dexterous capability. It discusses the basic principles and the challenges faced of the dexterous manipulation, and also suggests improvements needed in current foundation models for manipulation to enable general manipulation to achieve dexterity.
Bicchi (2000) offers a thorough and widely accepted definition: dexterous manipulation is the capability of changing the position and orientation of the manipulated object from a given reference configuration to a different one, arbitrarily chosen within the hand workspace. Based on this definition, the dexterous manipulation can be described as: based on the designed end-effector, determining a sequence of contact points and the forces/torques to be exerted on the object, and control the whole-body to accomplish a specific task.
Based on this definition, the challenges of dexterous manipulation lie in the design of the end-effector, determining a sequence of contact points and forces/torques, and whole-body control. The process of determining a sequence of contact points and forces/torques can be divided into model-based approach and the learning-based approach. The model-based approach is interpretable, explicitly showing the factors to consider in dexterous manipulation. Therefore, this section explains both model-based approach and learning-based approach.
9.5.1 End-effector Design.
Currently, there are two primary approaches to designing end-effector. The first approach customizes the end-effector for specific tasks. The second approach makes the multi-fingered end-effector resemble a human hand. The end-effector designed with the first approach is usually easier to control because it has fewer degrees of freedom compared to the end-effector designed with the second approach. In Billard and Kragic (2019), dexterity is divided into two types: extrinsic dexterity and intrinsic dexterity. Extrinsic dexterity involves using external support, such as friction, gravity, and contact surfaces, to compensate for the lack of degrees of freedom. Intrinsic dexterity refers to the hand’s ability to manipulate objects using its own degrees of freedom. Therefore, the first approach still has certain limitations for general manipulation. The design of the end-effector is related to robot proprioception $S_{robot}$ and this survey includes this topic in the State module.
The first approach requires manual design, extensive testing, and continual adjustments. In Stella et al. (2023), LLMs are used for designing end-effector. However, this area is still in its early exploration stages. Using LLMs for end-effector design generates text descriptions, which still need to be manually translated into designs. This process is not fully automated. If we could develop modules for rotational and translational joints, and use something like protein structure prediction networks (Jumper et al. (2021)), training a foundation model to output graph including these joints could help reduce the challenges of manual design. As for the second approach, the human hand has many sensors and actuators. This makes it nearly impossible to design a robotic hand that closely resembles the human hand. Therefore, it’s essential to design the sensors and actuators carefully.
9.5.2 Model-Based Generation of a Sequence of Contact Points and Forces/Torques
When solving simple tasks, contact points can remain fixed. However, for complex tasks, contact points need to change. Thus, a sequence of contact points and forces/torques is required, achieved through regrasping or finger gaiting. The sequence of contact points and forces/torques are positively correlated with the trajectory of the motion and wrench of the manipulated object. When the wrench and motion of the manipulated object at a given moment are determined, they can be mapped to the corresponding contact points and forces/torques between the end-effector and the manipulated object. Therefore, we will explain the pipeline of this mapping relationship for a specific moment.
Assuming the motion trajectory of the object is already obtained, the target wrench can be calculated based on the object’s mass and inertia. Contact points can be optimized based on object geometry, object material, and end-effector geometry etc, using appropriate metrics (Ferrari et al. (1992)). Alternatively, they can be generated using a knowledge-based approach (Stansfield (1991)). Subsequently, a coordinate system is established at the centroid of the object. Based on the location of the contact points, the target wrench is converted into fingertip force. Finally, the fingertip force is converted into the forces/torques required by the actuators through inverse kinematics (Okamura et al. (2000)). The pipeline contact points and forces/torques mentioned above are derived sequentially. In practical applications, both can be optimized simultaneously (Xu et al. (2024)).
Rolling and sliding at the contact points can hinder task execution. Tasks like pushing and pulling don’t need strong stability between the end-effector and the object. However, for tasks like grasping, stability is crucial and the system must keep stable contact even when disturbed. Form-closure and force-closure are common approaches. Form-closure depends on the end-effector’s design. Force-closure involves using force control to counteract external forces, like gravity. Current grasping methods mainly increase grip force to boost friction and achieve force-closure. However, to achieve precise wrench control, it is essential to understand the coefficient of friction represented by the object’s material. Tactile-based slip prediction can also enhance stability (Veiga et al. (2015)), but it requires high real-time performance.
From the above analysis, it is evident that choosing appropriate contact points and forces/torques requires considering the object’s geometry, mass, inertia, material, and friction parameters, as well as the end-effector’s geometry, material, and actuator capability. Tactile-based slip prediction is also needed to ensure the stability of the contact points. The end-effector’s properties are known in advance, but the object is unseen. Therefore, estimating the object’s properties should mimic human perception, using vision, sound, and touch to estimate the object’s properties. Repeated interactions can then refine the accuracy of these estimations.
9.5.3 Learning-Based Generation of a Sequence of Contact Points and Forces/Torques
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the robot.
Hand motion capture can be categorized into camera-based, IMU-based, and electromagnet-based methods. Using camera-based methods, human motion is estimated through pose estimation and retargeted to a multi-fingered end-effector. However, the accuracy of pose estimation needs improvement and occlusions can cause information loss. IMU-based methods are typically used to estimate wrist pose but miss fine-grained finger motion. Additionally, IMU used for pose estimation are prone to drifting over time. Electromagnet-based methods primarily use electromagnetic field gloves to estimate finger motion, but they lack information about wrist pose. At the same time, IMU-based methods and electromagnet-based methods both lack the vision data needed for policy training. DEXCAP (Wang et al. (2024)) designs a portable hand motion capture system. It uses Rokoko gloves to capture finger motion. A Realsense T265 camera is added to the wrist of the gloves to monitor wrist pose. Additionally, a Realsense L515 LiDAR camera is mounted on the chest to sense the environment. However, the glove’s design lacks tactile feedback. Therefore, to collect data for a multi-fingered end-effector, hardware improvements are necessary.
Two main learning-based methods for dexterous manipulation are imitation learning (Ze et al. (2024)) and reinforcement learning (Ma et al. (2023b)). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (Chi et al. (2023)) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. However, it remains uncertain whether diffusion policy is suitable for one-model to multiple tasks approach. It can also use an existing VLM for fine-tuning (in Sec. 7). Fine-tuning with a VLM allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (Brohan et al. (2023)). Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (Ye et al. (2023a)) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (Gupta et al. (2021)) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
Current learning methods each have their strengths and weaknesses. Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the ”Model-Based Generation of a Sequence of Contact Points and Forces/Torques”. ManiFoundation(Xu et al. (2024)) generates contact point and force heatmaps based on task goals, object material, object geometry and end-effector geometry. However, ManiFoundation lacks sensors such as tactile for real-time servoing of the manipulated object’s motion.
9.5.4 Whole-body Control
The above discussion primarily focuses on the contact between the end-effector and the object. However, whole-body control is still needed in dexterous manipulation. For example, in a polishing robot, force-position hybrid control of the robotic arm is often required to manage the trajectory of contact points and forces/torques. Mobile manipulation is essential for dexterous manipulation reachability. This idea is inspired by how humans handle objects. For example, when playing badminton, people use their waists, shoulders, elbows, and wrists together to hit the shuttlecock further. This aspect is often overlooked by current foundational models for manipulation. Although LEO (Huang et al. (2023a)) can provide poses for both navigation and manipulation, it still does not address the synchronization issue between the two.
For whole-body control, the focus is on low-level control issues. A straightforward idea is to expand the action space of the policy module’s robot foundation models to include all joints of the robot. However, as the output dimensions increase, end-to-end training methods are more likely to diverge. Therefore, most current models output cartesian space poses and force/torques. These outputs are then optimized and converted into position or torque for each joint through a post-processing module (Haviland and Corke (2021)). To address end-to-end whole-body control issues, principal research is needed to facilitate network training and deployment.
@@ 9.6 How to establish a benchmark?
Current research on foundation models for manipulation focuses on various tasks, including interaction, hierarchical tasks, perception, detecting pre- and post-conditions, policy, and manipulation data generation. Therefore, a benchmark for foundation models for manipulation should include a comprehensive framework with diverse tasks. This framework should test individual tasks and tasks that involve connecting different modules. Building a benchmark should include existing datasets. However, since different simulators have unique physics engines and renderers, the benchmark should include various simulators and separately test all existing datasets in different simulators.
Table. 1 lists the benchmarks used in current RFMs, highlighting a lack of standardization. This inconsistency hinders the development of RFMs for three main reasons. Firstly, current RFMs are tied to the specific parameters of each robot, such as the choice of sensors, camera pose, and the robot’s degrees of freedom. These factors prevent RFMs from being easily transferred across different robots. Secondly, testing the generalization and success rate of general manipulation capability requires a wide range of scenes and tasks. Thirdly, there is no standardized metric for assessing general manipulation capability.
As for the RFMs are not transferable between different robots. The issue arises from focusing solely on testing RFM algorithms without considering hardware, which is an ineffective approach. General manipulation requires whole-body control. Thus, evaluating the generalization and success rate of RFMs should involve both algorithms and hardware, unlike in computer vision where only algorithms are considered. To address this, the simulation benchmark should include an easy interface for importing various robot hardware configurations.
As for the requirment of a wide range of scenes and tasks. Although iGibson (Li et al. (2021)) and BEHAVIOR-1K (Li et al. (2023a)) support simulating a variety of household tasks with high realism, they are still manually constructed. In Section 8, we discuss how foundation models can automate the generation of scenes and tasks. Using foundation models to create numerous scenes, combined with VLMs for accuracy checking and minimal human intervention, could be a valuable approach to explore.
As for the metric for assessing general manipulation. The current evaluation standards mainly focus on success rates. However, in real-world applications, user preferences should also be considered. For instance, the system’s real-time performance is important. Most algorithms focus on building the generalization of skills. They often overlook the amount of data and speed required for RFMs to learn a new skill. Therefore, evaluation metric should also include the learning ability of RFMs.
Overall, to assess the ability for general manipulation, methods used for testing medical robots can be referenced. Start with extensive testing in simulation environments, followed by limited tests in real-world settings. Continue evaluating the general manipulation capability during the product’s application phase.
@@ 10 Conclusion
The impressive performance of foundation models in the fields of computer vision and natural language suggests the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, current research lacks consideration of a general manipulation framework. Thus, this paper proposes a general manipulation framework based on the development of robot learning for manipulation and the definition of general manipulation. It also describes the opportunities that foundation models bring to each module of the framework.
We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0), the current research has a high probability of achieving L0.
Then, we discuss the following points: (1) the logic and implementation strategies of the designed framework, (2) how to design the modules of post-conditions detection and post-hoc correction, due to limited research on foundation models for these areas, (3) the learning capability required for general manipulation, (4) the strengths and weaknesses of current RFMs, (5) how to make the general manipulation with the dexterous capability, and (6) how to establish a foundation models for manipulation benchmark.
Additionally, the proposed framework has certain limitations: (1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates and safety for specific tasks. Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework is limited to static environments, rigid objects, short-horizon tasks, improving old skills, and low precision requirements for contact points and forces/torques. Further research is needed to advance from L0 to L5.
Finally, foundation models present opportunities for each module of the framework, but many challenges still remain:
@@
    Interaction Human interaction involves not only language but also gestures and actions. Incorporating multimodal inputs into interaction modules can enhance recognition capability.
@@
    Hierarchical of skills The hierarchy of skills still has many unconsidered factors, such as achieving tasks in the shortest time with the highest efficiency, and how to generate strategies for dynamic scenes.
@@
    Pre- and post-conditions detection Current research on post-condition detection primarily focuses on detection after robot execution. However, this delay is unacceptable. Therefore, it is necessary to implement failure detection and analysis of failure reasons during the robot execution.
@@
    State The representation of state requires integration of multiple modalities, such as touch and hearing. Additionally, it’s important to consider the opportunities that foundation models can bring to active perception.
@@
    Policy Current research on RFMs primarily involves fine-tuning VLMs. This approach deprives RFMs of the ability to self-explore. The extensive parameters of RFMs require significant computational resources for training and real-time reference, and model training also needs abundant data. Additionally, there is a lack of a unified benchmark for evaluating different RFMs.
@@
    Environment Transition Module The foundation models inherently contain abundant physical priors. Applying foundation models to build a highly realistic physical model assist reinforcement learning training is a direction worth exploring.
@@
    Data Generation The accuracy of data generated by LLMs and VGMs remains insufficient, necessitating appropriate check module and data cleaning algorithms.